OpenStudy (anonymous):

A bicycle is marked 40% off the original price of $150. It is then taxed at 7.5%. What is the final total cost of the bicycle? $101.25 $97.50 $64.50 $96.75

jimthompson5910 (jim_thompson5910):

What is 40% of $150?

OpenStudy (anonymous):

idk.. i feel so dumb :'(

jimthompson5910 (jim_thompson5910):

its ok, you're not dumb 40% of 150 = 0.4 times 150 = ???

jimthompson5910 (jim_thompson5910):

its ok to use a calculator if you don't know it off the top of your head

OpenStudy (anonymous):

60?

jimthompson5910 (jim_thompson5910):

so applying the 40% discount means you save $60 and you instead pay 150 - 60 = 90 dollars this is only true if there are no taxes but there are taxes so you need to add on 7.5% of 90 dollars to 90 dollars Price without taxes = $90 Tax = 7.5% of 90 = 0.075*90 = 6.75 Price with taxes = Price without taxes + Tax Price with taxes = 90 + 6.75 Price with taxes = 96.75

OpenStudy (anonymous):

wow thanks you're so smart!

jimthompson5910 (jim_thompson5910):

you're welcome